OpenStudy (anonymous):

Write 1/5 + 1/6 + 1/7 in sigma notation

jimthompson5910 (jim_thompson5910):

What's changing from term to term?

OpenStudy (anonymous):

1/5?

jimthompson5910 (jim_thompson5910):

how are 1/5 and 1/6 different

OpenStudy (anonymous):

\(\Large \sum _{ n=5 }^{ 3 }{ \frac { 1 }{ n } } \)

OpenStudy (anonymous):

\[\sum_{n=5}^{7}\frac{ 1 }{ n }\]